completing squares
directions: find the value of c that makes the expression a perfect square trinomial. Then write the expression as the square of a binomial.
question: x2 + 10x + c


my work: 10
-- = 5 then (5)2= 25 then x2 + 10x + 25 = 0 then (x + 5) 2 = 0
2


then I square rooted both sides which = x + 5 = 0 x = + 5
-5 -5 -
-----------
Well I don't know if I did it right someone please help.

Answer by stanbon(48569)

write the expression as the square of a binomial.
question: x^2 + 10x + c
--------------
x^2 + 10x + ? = -c + ?
x^2 + 10x + (10/2)^2 = -c + (10/2)^2
x^2 + 10x + 25 = -c + 25
(x+5)^2 = 25-c
==================
Cheers,
Stan H.